# A Survey on Integrated Sensing, Communication, and Computation

### v1

@@ III-B2 Radar Architectures
In this part, five radar architectures are introduced, as shown in Fig. 4.
Figure 4: Illustration of five radar architectures.
According to the locations and numbers of transmitters and receivers, radar architectures are categorized into mono-static, bi-static, and multi-static types with each having its pros and cons [56, 112, 57, 113], which are discussed as follows.
  • Mono-static Radar: The mono-static radar refers to that the transmitter and receiver are co-located [56, 112]. As a result, this architecture enjoys a concise structure and can be deployed conveniently. More importantly, signal processing is simplified as the synchronization of transmit and receive signals is easy. However, due to the simple structure, its radar cross section (RCS) is limited, leading to a relatively poor performance. Besides, continuous waveforms are not feasible in mono-static radar as they suffer from severe self-interference.
  • Bi-static Radar: In a bi-static radar system, one transmitter and one receiver are separately deployed [57, 113]. Such systems are featured by the transmitter-target-receiver triangulation. Under this setting, no self-interference is caused at the receiver, leading to a more flexible choice of waveforms. The transmit and receive antennas can be implemented and optimized independently and the RCS-like range and angular resolutions can be improved via spatial filtering. However, the separate design brings more difficult time synchronization between the transmitter and the receiver. Additional information like locations and waveform types should be shared, leading to a higher cost for maintenance.
@@
To improve radar’s sensing ability for object detection in a wider area, early radar systems scan the target via the rotation of a single antenna [114]. In this system, an antenna periodically emits a probing signal and receives the reflected signal (i.e., the same antenna is often used for both transmitting and receiving). As a result, time-division duplex should be carefully designed and waveforms like CW and FMCW cannot be applied. Benefiting from the advancement of multi-antenna techniques, new solutions are proposed for improving the detection coverage of radar, as described below.
  • Phased Array Radar: Phased array radar comprises multiple antenna (or radiating) elements arranged in a particular order on a curved or plane surface. Its transmit waveform beam can be electronically steered to points in different directions and distances via controlling the power and phases among these antennas without changing their locations [58, 115, 116]. For example, by letting two radiators transmit a signal in the same phase shift, a waveform superposition is achieved and the signal strength is enhanced in the main direction and attenuated in other directions. By carefully designing the power and phase control strategy, phased array radar enjoys the advantages of high antenna gain with great side-lobe attenuation, arbitrary directions scanning, high beam agility, multi-functional operations via simultaneous generation of multiple beams, and enhanced robustness, which allows failure of some components. However, the complicated structure that needs to simultaneously control all antennas leads to higher costs and problems like low-frequency agility, deformation of the antenna pattern during beam steering, and limited scanning range.
  • Multiple-input and Multiple-output (MIMO) Radar: In MIMO radar systems, multiple antennas are equipped at both the transmitter and the receiver [59, 117, 118]. In such a system, each transmit antenna radiates arbitrary waveforms independent of other transmit antennas. At the receiver, the echo signal can be mapped with a single transmit antenna due to the independence of waveforms. By paying the cost of higher signal processing complexity, MIMO radar can design more adaptive transmit/receive beamforming and more flexible probing signals to achieve higher detection resolution and better parameter identifiability [119, 120, 121].
@@ III-B3 Intelligent Radar Sensing Systems and Applications
Radar sensing shows its superiority compared with other sensing techniques in many aspects, including long detection distance, high penetration property, easy to be equipped in vehicles. As a result, it is widely used in many military applications including military target detection, tracking, and recognition, and civil applications such as disaster assessment, land use and classification, and weather detection [62]. However, as the demands of empowering many intelligent services like auto-driving, human detection, and activity classification, modern radar systems encounter new requirements in terms of high accuracy, robustness, and real-time ability. This is beyond the capability of traditional radar signal processing techniques. To tackle this problem, AI, which has shown powerful abilities in many other areas like signal processing in biomedical and communication systems (see, e.g., [122, 123]), is applied for enabling intelligent radar signal processing.
One trend for intelligent radar signal processing aims at recognizing the radar radiation sources by processing the received waveforms with AI models [60], including two application scenarios. One is to distinguish different radar types, e.g., military radar or civil radar [61]. The other is to identify individual radar emitter parameters including the radiation characteristic, qualification of components, and so on [124]. The data used for AI model training includes pulse description words [125], amplitude-phase shift [126], I/Q time sequences [127], autocorrelation features [128], etc. Generally, to enhance the training performance, these raw data are pre-processed via e.g., denoising [129], missing data processing, and features extraction [127].
More literature about intelligent radar signal processing directly utilizes AI models to process the radar image, which refers to the echo wave power distribution in the delay-Doppler domain, for target detection, recognition, and natural scene analysis [62]. For example, AI models are used for canceling the interference and noise in synthetic aperture radar (SAR) images (see, e.g., [130]) and target detection or recognition (see, e.g.,[131, 132]). Particularly, adopting machine learning algorithms for processing micro-Doppler and range-Doppler images is widely used in many intelligent applications such as unmanned aerial vehicle (UAV) detection [133], human activity classification [134], and gait recognition [135].
@@ III-C Multi-modal Sensing Technology
A sensory modality refers to the way one entity is observed or perceived, e.g., via radar, camera, microphone, etc. Compared with unimodal sensing like radar sensing, multi-modal sensing can collect more information to achieve better task performance, which is mathematically proved in [30]. Besides, the common information can be obtained by different modalities, making multi-modal sensory data more robust to noise and sensory deficiency, i.e., data of one modality can provide complementary or supplementary information to others [31]. For example, in an audio-visual speech recognition task, the visual information helps enhance the speech recognition performance when the audio signal is noisy [31]. Moreover, the success of the recent deep generative models enables creating one modality of sensory data from another [111], such as using a text to generate a high-resolution image.
To enjoy the above advantages and based on the fact that data in the real world have multiple modalities, multi-modal sensory data has been the fuel to empower AI in many applications, including scene application, action recognition, video transcription, image captioning, and so on [31, 111, 136]. On the other hand, data of different modalities is heterogeneous in the forms (e.g., voice, image, and text) and the noise levels. The relation between different modalities is complex. Machine learning algorithms are currently the most powerful and widely adopted methods for multi-modal signal processing. These give rise to a new research area namely multi-modal machine learning. Subsequently, the key issues of multi-modal machine learning are discussed, followed by the introduction of the applications of multi-modal sensing.
@@ III-C1 Key Issues in Multi-modal Machine Learning
  • Representation: Representation is the information distilled from the data that intelligent systems can use [137]. Generally, representation and feature are interchangeably in AI tasks [31]. Many efforts have been put into efficiently extracting useful representations for AI tasks based on unimodal data, e.g., representation learning [137], feature detection [138], convolutional neural networks based feature extraction [139]. However, representation extraction from multi-modal sensory data faces new challenges. On one hand, data in each modality is obtained from different types of sensors with heterogeneous forms and noise levels. On the other hand, in many scenarios, data in some modalities is missing. The authors in [66] pointed out that a good multi-modal representation should satisfy the following novel requirements: 1) Similarity in representation should lead to the same similarity in the corresponding results; 2) Multi-modal representation should be readily obtained even in the absence of certain modalities; 3) Missing modalities can be generated using other observed ones. As a result, it remains a challenging problem to construct a good representation for multi-modal data.
    The current methods for multi-modal representation can be categorized into two types, i.e., joint representation and coordinated representation [31, 32]. Joint representation projects data of all modalities into the same latent representation space, which is the most adopted method [31]. As well as using the conventional neural networks, autoencoders are utilized for pre-training the unlabeled data (e.g., [67, 140, 141]) and probabilistic graphical models like deep Boltzmann machines are adopted to deal with the issue of missing modality via using their generative nature [66]. In coordinated representation, different modalities have their own representations but are coordinated via setting a constraint of similarity (e.g., Euclidean distance) or structure (e.g., partial order) [142].
  • Translation: Multi-modal translation refers to the ability of modality transfer i.e., generating some modalities of one entity with one source modality [31]. There are two kinds of translation approaches, i.e., example-based and generative approaches.
    • Example-based approach: In this approach, a dictionary is constructed using the training data samples. For each entity, the source modality and target modality of one entity are stored in the dictionary with one-to-one correspondence [31]. The simplest way called the retrieval-based model for modality translation is finding the closest entry in the dictionary and using it as the translation result [63]. Another way, called the combination-based model, takes one step further. It combines multiple entries in the dictionary with a designated way to create a better translation result [31]. For example, the authors in [143] find the most similar $k$ items of one source image and utilize their captions to generate a target descriptive sentence of the source image. Although the idea of the example-based approach is simple, it calls for expensive costs for the construction and inference of the dictionary especially when the number of items is large [31].
@@
    • Generative approach: In this approach, generative AI models are trained for modality translation [31]. Specifically, two kinds of methods are included. One is called the grammar-based model, which first extracts high-level concepts from the source modality. Then, these concepts are integrated via a pre-defined grammar-based generation procedure for creating a target modality [144]. By using predefined templates and restricted grammars, this method can generate syntactically or logically correct target instances. However, this method prefers creating formulaic rather than creative translations, and a separate model and training dataset are required to detect each concept, leading to complex pipelines for concept detection [31]. The other method is the encoder-decoder model. It first encodes a source modality into a vectorial representation and then utilizes a decoder to generate the target modality [145, 146]. Although this method calls for large quantities of data for training, it is widely used in many generative tasks, like text, image and sound generation, image caption, and video description.
@@
  • Alignment: Multi-modal alignment is defined as finding the relations and correspondences among several modalities of one entity [31]. With good alignment, the performance of modality translation can also be enhanced. Multi-modal alignment can be divided into two categories, i.e., explicit alignment and implicit alignment. The former directly aligns the sub-components among different modalities, e.g., aligning a movie to a script. The latter latently aligns data of different modalities during AI model training, targeting enhancing the performance of many intelligent tasks, like speech recognition, machine translation, etc. There are three main issues to be addressed for multi-modal alignment. The first is that not all elements in one modality have correspondences in another, especially in the case of missing modalities. The second is the difficulty in designing a similarity metric for different modalities. The third is the scarce datasets with explicitly annotated alignments [31]. In the sequel, the methods for explicit and implicit alignment are introduced.
@@
    • Explicit alignment methods: For the case where there are no alignment labels, i.e., labeled correspondences between two modalities, unsupervised explicit alignment algorithms are required. One basic approach is called dynamic time warping (DTW) [64]. It measures the similarity between the time sequences of two modalities by time warping and aims at finding an optimal match between them. The alignment performance of such an approach highly depends on the similarity metric. Then, an extended version of DTW is proposed by introducing canonical correlation analysis (CCA) [147, 148]. As a result, both alignment (through DTW) and learning the mapping (through CCA) between the sequences of different modalities are enabled. The DTW approach and its extension align multi-modal data via linear transformations. To deal with the non-linear relations, a deep canonical time-warping approach is proposed in [149]. To further learn the complex alignment relations between modalities, various kinds of graphical models are also popular for multi-modal data alignment, including generative graphical models [150], factored hidden Markov models [151], dynamic Bayesian networks [152], etc. For the case where there are alignment labels, supervised explicit alignment algorithms are proposed, most of which take inspiration from the unsupervised methods mentioned above, in addition to utilizing the alignment labels to train a better similarity measure (e.g., [153, 154, 155]).
@@
    • Implicit alignment methods: As mentioned, implicit alignment methods try to latently align multi-modal data for completing a machine learning task instead of directly aligning the instances of multiple modalities. The early methods adopted graphical models [156], where the mapping between modalities should be manually constructed. Other methods directly use the neural networks for implicit alignment. Specifically, various kinds of different attention models have been proposed for better performance in question answering and better model interpretability [157], including hierarchical [158], stacked [159], and episodic memory attention [160].
@@
  • Fusion: Multi-modal fusion refers to integrating information from multiple modalities of the same phenomenon for completing a prediction task, such as classification and regression [31]. As mentioned, using multi-modal information for task completion leads to more robust performance due to their complementary property. However, there are two challenges [31]. On one hand, the signals of different modalities may not be temporally synchronized and exhibit different noise levels. On the other hand, it is difficult to build up models that can explicitly exploit the complementary properties of multiple modalities. To address the above challenges, three types of multi-modal fusion approaches are proposed, i.e., early, late, and hybrid fusion [65]. Early fusion refers to integrating the representations or features of multi-modal data and the integrated representations or features are used for completing the downstream task. It can learn the correlation and interactions of different modalities in the early stage and only a single model is trained. Late fusion is similar to ensemble learning. It fuses the decisions based on unimodal data using mechanisms like averaging [161], voting schemes [162], or a learned model [163, 164]. Late fusion uses one predictor for each modality, allowing more flexibility and is more robust to modality missing. Hybrid fusion, as its name suggests, combines the above two types to enjoy the advantages of both.
  • Co-learning: Co-learning refers to exploiting knowledge from a (resource-rich) modality to aid the modeling of another (resource-poor) modality when the latter is lack of data, polluted, or has unreliable labels [31]. To achieve high co-learning performance, several challenges should be overcome. First, the learned representation of the resource-rich modality may be task-relevant, leading to a limited performance when transferring its knowledge across different modalities in different tasks. Second, the instance numbers and labels are unbalanced. Moreover, the noise levels and the reliability of the labels are heterogeneous. According to whether the multi-modal instances of the same entity are aligned during training, the co-learning approaches are categorized into three types [31].
    • Aligned instances: In this case, two methods are adopted. One is co-training, where instances of two modalities bootstrap each other with labels for unlabelled instances [31]. This method, however, may lead to unbiased training samples and cause overfitting. The other method uses transfer learning-based multi-modal representation learning, which transfers the representation of one modality to another (e.g., [66]).
@@
    • Unaligned instances: In this case, the training instances of different modalities are not necessary from the same entities. The first method uses transfer learning to achieve coordinated multimodal representations, e.g., using text messages to improve visual representations for image classification [165]. The second is conceptual grounding. It learns semantic meanings or concepts based on data of multiple modalities, including language, vision, and sound, by finding a common latent space of all multi-modal representations (e.g., [166]). The last one is zero-shot learning, which recognizes a concept without explicitly seeing any examples of it [31]. For example, it can use a neural network to map the text-based message with the extracted visual features.
    • Hybrid case: In this case, approaches use a pivot modality to learn coordinated multi-modal representations using e.g., bridge correlational neural network, or rely on the existence of large datasets from a similar or related task [31].

### v2

@@ III-B2 Radar Architectures
In this part, five radar architectures are introduced. According to the locations and numbers of transmitters and receivers, radar architectures are categorized into mono-static, bi-static, and multi-static types with each having its pros and cons [83, 84], which are discussed as follows.
  • Mono-static Radar: The mono-static radar refers to that the transmitter and receiver are co-located [83]. As a result, this architecture enjoys a concise structure and can be deployed conveniently. More importantly, signal processing is simplified as the synchronization of transmit and receive signals is easy. However, due to the simple structure, its radar cross section (RCS) is limited, leading to a relatively poor performance. Besides, continuous waveforms are not feasible in mono-static radar as they suffer from severe self-interference.
  • Bi-static Radar: In a bi-static radar system, one transmitter and one receiver are separately deployed [84]. Such systems are featured by the transmitter-target-receiver triangulation. Under this setting, no self-interference is caused at the receiver, leading to a more flexible choice of waveforms. The transmit and receive antennas can be implemented and optimized independently and the RCS-like range and angular resolutions can be improved via spatial filtering. However, the separate design brings more difficult time synchronization between the transmitter and the receiver. Additional information like locations and waveform types should be shared, leading to a higher cost for maintenance.
@@
To improve radar’s sensing ability for object detection in a wider area, early radar systems scan the target via the rotation of a single antenna [85]. In this system, an antenna periodically emits a probing signal and receives the reflected signal. As a result, time-division duplex should be carefully designed and waveforms like CW and FMCW cannot be applied. Benefiting from the advancement of multi-antenna techniques, new solutions, i.e., phased array radar and multi-input and multi-output (MIMO) radar as shown in Fig. 5, are proposed for improving the detection coverage of radar, as described below.
Figure 5: Illustration of phased array radar and MIMO radar.
  • Phased Array Radar: Phased array radar comprises multiple antenna (or radiating) elements arranged in a particular order on a curved or plane surface. Its transmit waveform beam can be electronically steered to points in different directions and distances via controlling the power and phases among these antennas without changing their locations [86, 87]. By carefully designing the power and phase control strategy, phased array radar enjoys the advantages of high antenna gain with great side-lobe attenuation, arbitrary directions scanning, high beam agility, multi-functional operations via simultaneous generation of multiple beams, and enhanced robustness, which allows failure of some components. However, the complicated structure that needs to simultaneously control all antennas leads to higher costs and problems like low-frequency agility, deformation of the antenna pattern during beam steering, and limited scanning range.
  • MIMO Radar: In MIMO radar systems, multiple antennas are equipped at both the transmitter and the receiver [88, 89]. In such a system, each transmit antenna radiates arbitrary waveforms independent of other transmit antennas. At the receiver, the echo signal can be mapped with a single transmit antenna due to the independence of waveforms. By paying the cost of higher signal processing complexity, MIMO radar can design more adaptive transmit/receive beamforming and more flexible probing signals to achieve higher detection resolution and better parameter identifiability [90].
@@ III-B3 Intelligent Radar Sensing Systems and Applications
Radar sensing shows its superiority compared with other sensing techniques in many aspects, including long-detection distance, high penetration properties, and ease to be equipped in vehicles. As a result, it is widely used in many military applications including military target detection, tracking, and recognition, and civil applications such as disaster assessment, land use and classification, and weather detection [91]. However, as the demands of empowering many intelligent services like auto-driving, human detection, and activity classification, modern radar systems encounter new requirements in terms of high accuracy, robustness, and real-time ability. This is beyond the capability of traditional radar signal processing techniques. To tackle this problem, AI, which has shown powerful abilities in many other areas like signal processing in biomedical and communication systems (see, e.g., [92, 93]), is applied to enable intelligent radar signal processing. According to [94], such techniques can be categorized into five types, i.e., intelligent waveform design, intelligent array design for MIMO radar, radiation source recognition, target detection and recognition, and interference recognition and suppression, which are elaborated below.
Intelligent waveform design refers to using AI models to design waveforms for interference suppression and main lobe gain enhancement of MIMO radar. To suppress the interference of radar signals from other systems (e.g., wireless communication systems) when spectrum sharing is allowed, a reinforcement learning strategy based on deep deterministic policy gradient (DDPG) is adopted to select phases for a phase-coded waveform in [95]. A dual Q-learning algorithm with long short-term memory (LSTM) is designed to enable radar signals to use idle sub-bands in [96]. On the other hand, to enhance the main lobe gain and minimize the sidelobe radiation in MIMO radar, deep neural networks (DNNs) such as deep residual neural networks [97] and feed-forward neural networks [98], are adopted.
Intelligent array design for MIMO radar refers to partitioning the antenna array into multiple subarrays using machine learning methods. The waveforms of different subarrays are designed to be orthogonal in order to achieve a diversity gain [99, 100]. In particular, a convolutional neural network is designed in [101] to adaptively select the optimal subarray in a time-varying environment. By training a lightweight convolutional neural network generated by phased-MIMO array manifold matrix, a sparse transmit subarray is selected using a proposed interleaved partitioning method to transmit the reduced-dimensional radar data in [102].
Radiation source recognition identifies and classifies various radar and communication signals based on their emitted waveforms [103]. Its main goals include detecting and classifying low probability of interception (LPI) radar signals, enhancing the detection accuracy of passive radar systems, and resisting interference or adversarial attacks. Specifically, convolutional and recurrent neural networks are widely used for LPI signal classification [104, 105]. For improving passive radar accuracy, feature extraction methods like the Wigner Ville Distribution (WVD) are utilized to estimate direct-path signals [106]. Moreover, adversarial training and detection strategies are developed to address vulnerabilities to adversarial attacks, which can disrupt signal classification performance [107].
Target detection and recognition involve identifying and classifying objects such as vehicles, ships, and aircraft based on radar signals [91]. In this case, the features such as high-range resolution profiles, Doppler signatures, and synthetic aperture radar images are extracted and fed into different kinds of AI models, including convolutional neural networks, stacked autoencoders, and LSTM networks for fulfilling the tasks [108]. To deal with limited training data, techniques such as generative adversarial networks [109] and transfer learning [110] are employed to expand the available data set and boost recognition accuracy. Additionally, semi-supervised [111] and unsupervised learning [112] methods are used when labeled data is scarce, improving recognition performance even with minimal human-labeled input. Moreover, adversarial training techniques are developed to protect the models from adversarial attacks that aim to mislead the classifier, ensuring robust performance in challenging environments [113].
AI has been a powerful tool for jamming/interference recognition and suppression in radar systems. In [114], a fractional Fourier transform based multifeature fusion network uses the local and global features of interference signals and the attention mechanism to improve the interference recognition performance. A multi-scale attention network (MANet) can extract different features through multi-scale dilated convolution, realizing the classification and recognition of target signals and interference signals in ultra-wideband (UWB) signals [115]. What’s more, by using random signal enhancement and unsupervised domain adaptation to mitigate environmental dependencies in each aspect, a clutter cancellation approach is proposed in [116] under the varying environment.
@@ III-C Multi-modal Sensing Technology
A sensory modality refers to the way one entity is observed or perceived, e.g., via radar, camera, microphone, etc. Compared with unimodal sensing like radar sensing, multi-modal sensing can observe or perceive entities through multiple sensing modalities to achieve better task performance. Particularly, the authors in [31] mathematically proved that the fusion of data from multiple modalities increases the amount of statistical information and improves the accuracy of learning tasks. Besides, common information can be obtained by different modalities, making multi-modal sensory data more robust to noise and sensory deficiency, i.e., data of one modality can provide complementary or supplementary information to others [32]. For example, in an audio-visual speech recognition task, the visual information helps enhance the speech recognition performance when the audio signal is noisy [32]. Moreover, the success of the recent deep generative models enables creating one modality of sensory data from another [81], such as using text to generate a high-resolution image or a video. This development can bridge the gap between different types of perceptual data, and its applications in computer vision, natural language processing, and other fields are more diverse and creative.
The key design principle for multi-modal sensory data processing is to enhance the performance of downstream tasks by effectively transforming the original data into latent representations suitable for tasks such as object detection [117]. This process concerns both efficient encoding and the fusion of common information from diverse sources to maximize the relevance of shared features. The challenge lies in managing the inherent heterogeneity of the data streams, where the variations in quality, format, and temporal alignment must be accounted for to ensure seamless integration [118]. The processing procedures of data in different modalities vary and the relation among them is complex. To overcome these challenges and enjoy the previously mentioned benefits of multi-modal sensing, machine learning algorithms are currently the most powerful and widely adopted methods for multi-modal signal processing [32, 81, 119]. These give rise to a new research area namely multi-modal machine learning. Subsequently, the key issues of multi-modal machine learning are discussed, followed by the introduction of the applications of multi-modal sensing.
@@ III-C1 Key Issues in Multi-modal Machine Learning
  • Representation: Representation or called feature is the information distilled from the data that intelligent systems can use [120]. Many efforts have been put into efficiently extracting useful representations for AI tasks based on unimodal data, e.g., representation learning, feature detection, and convolutional neural networks based feature extraction. However, representation extraction from multi-modal sensory data faces new challenges. On one hand, data in each modality is obtained from different types of sensors with heterogeneous forms and noise levels. On the other hand, in many scenarios, data in some modalities is missing. Moreover, it’s difficult to measure the correlation and learn a joint distribution of different modalities [121]. As pointed out in [122], a good multi-modal representation should satisfy the following requirements: 1) Similarity in representation should lead to the same similarity in the corresponding results; 2) Multi-modal representation should be readily obtained even in the absence of certain modalities; 3) Missing modalities can be generated using other observed ones. The current methods for multi-modal representation can be categorized into two types, i.e., joint representation and coordinated representation [32, 33]. Joint representation projects data of all modalities into the same latent representation space, which is the most adopted method [32]. As well as using the conventional neural networks, autoencoders are utilized for pre-training the unlabeled data (e.g., [123]) and probabilistic graphical models like deep Boltzmann machines are adopted to deal with the issue of missing modality via using their generative nature [122]. In coordinated representation, different modalities have their own representations but are coordinated via setting a constraint of similarity (e.g., Euclidean distance) or structure (e.g., partial order) [124].
@@
  • Translation: Multi-modal translation refers to the ability of modality transfer i.e., generating some modalities of one entity with one source modality [32]. How to effectively evaluate the performance of multimodal translation is a problem that needs to be solved. For example, although there is a single answer for tasks such as speech recognition, there is no correct answer for more tasks such as speech synthesis and media description, and there are multiple correct answers for language translation. There are two kinds of translation approaches, i.e., example-based and generative approaches. One is the example-based approach. In this approach, a dictionary is constructed using the training data samples. For each entity, the source modality and target modality of one entity are stored in the dictionary with one-to-one correspondence [32]. The simplest way called the retrieval-based model for modality translation is finding the closest entry in the dictionary and using it as the translation result [125]. Another way, called the combination-based model, takes one step further. It combines multiple entries (e.g., the most similar $k$ items) in the dictionary with a designated way to create a better translation result [32]. Although the example-based approach is simple, it calls for expensive costs for the construction and inference of the dictionary especially when the number of items is large [32]. The other is the generative approach. In this approach, generative AI models are trained for modality translation [32], including two methods. The first is called the grammar-based model, which first extracts high-level concepts from the source modality. Then, these concepts are integrated via a pre-defined grammar-based generation procedure for creating a target modality [126]. However, this method prefers creating formulaic rather than creative translations, and a separate model and training dataset are required to detect each concept, leading to complex pipelines for concept detection [32]. The second method is the encoder-decoder model. It first encodes a source modality into a vectorial representation and then utilizes a decoder to generate the target modality [127]. Although this method calls for large quantities of data for training, it is widely used in many generative tasks, like text, image and sound generation, image caption, and video description.
@@
  • Alignment: Multi-modal alignment is defined as finding the relations and correspondences among several modalities of one entity [32]. With good alignment, the performance of modality translation can also be enhanced. There are three main issues to be addressed for multi-modal alignment. The first is that not all elements in one modality have correspondences in another, especially in the case of missing modalities. The second is the difficulty in designing a similarity metric for different modalities. The third is the scarce datasets with explicitly annotated alignments [32]. To address these issues, two categories of techniques are proposed.
@@
    The first is the explicit alignment method. For the case where there are no alignment labels, i.e., labeled correspondences between two modalities, unsupervised explicit alignment algorithms are required. One approach is the linear transformation based approach called dynamic time warping (DTW) [128]. It measures the similarity between the time sequences of two modalities by time warping and aims at finding an optimal match between them. The alignment performance highly depends on the similarity metric. An extended version of DTW introduces canonical correlation analysis (CCA) to enable both alignment (through DTW) and learning the mapping (through CCA) between the sequences of different modalities [129]. To deal with the non-linear relations, the deep canonical time-warping approach [130], and various kinds of graphical models based approaches, e.g., generative graphical models [131], factored hidden Markov models [132], and dynamic Bayesian networks [133], are proposed. For the case where there are alignment labels, supervised explicit alignment algorithms are proposed, most of which take inspiration from the unsupervised methods mentioned above, in addition to utilizing the alignment labels to train a better similarity measure (e.g., [134, 135]).
@@
    The second is the implicit alignment method. It tries to latently align multi-modal data for completing a machine learning task instead of directly aligning the instances of multiple modalities. The early methods adopted graphical models [136], where the mapping between modalities should be manually constructed. Other methods directly use the neural networks for implicit alignment. Specifically, various kinds of different attention models have been proposed for better performance in question answering and better model interpretability [137], including hierarchical [138], stacked [139], and episodic memory attention [140].
@@
  • Fusion: Multi-modal fusion refers to integrating information from multiple modalities of the same phenomenon for completing a prediction task, such as classification and regression [32]. As mentioned, using multi-modal information for task completion leads to more robust performance due to their complementary property. However, there are two challenges [32]. On one hand, the signals of different modalities may not be temporally synchronized and exhibit different noise levels. On the other hand, it is difficult to build up models that can explicitly exploit the complementary properties of multiple modalities. To address the above challenges, three types of multi-modal fusion approaches are proposed, i.e., early, late, and hybrid fusion [141]. Early fusion refers to integrating the representations or features of multi-modal data and the integrated representations or features are used for completing the downstream task. It can learn the correlation and interactions of different modalities in the early stage and only a single model is trained. Early fusion is widely used in real-time processing, which is useful in scenarios where real-time integration of multi-modal data is required, such as in video surveillance or real-time sentiment analysis combining text and audio. Late fusion is similar to ensemble learning. It fuses the decisions based on unimodal data using mechanisms like averaging [142], voting schemes [143], or a learned model [144, 145]. Late fusion uses one predictor for each modality, allowing more flexibility and is more robust to modality missing. The application scope of late fusion is a complex decision-making which is suitable for tasks requiring complex decision-making based on multiple inputs, such as medical diagnosis combining images, patient records, and genetic data. Hybrid fusion, as its name suggests, combines the above two types to enjoy the advantages of both: early fusion for initial comprehensive feature representation and late fusion for refining these representations based on modality-specific processing. Besides, hybrid fusion methods are more flexible and adaptive to varying data conditions. They can handle scenarios where some modalities might be noisy, incomplete, or missing by allowing for modality-specific processing before integration. This adaptability is crucial in real-world applications where data quality and availability can vary significantly. One of the usage scenarios of hybrid fusion is the adaptive system that needs to adapt dynamically to changing conditions or varying data quality, like in environmental monitoring systems or smart cities.
@@
  • Co-learning: Co-learning refers to exploiting knowledge from a (resource-rich) modality to aid the modeling of another (resource-poor) modality when the latter is lack of data, polluted, or has unreliable labels [32]. To achieve high co-learning performance, several challenges should be overcome. First, the learned representation of the resource-rich modality may be task-relevant, leading to a limited performance when transferring its knowledge across different modalities in different tasks. Second, the instance numbers and labels are unbalanced. Moreover, the noise levels and the reliability of the labels are heterogeneous. According to whether the multi-modal instances of the same entity are aligned during training, the co-learning approaches are categorized into three types [32]. The first is the approch for aligned instances. In this case, two methods are adopted. One is co-training, where instances of two modalities bootstrap each other with labels for unlabelled instances [32]. This method, however, may lead to unbiased training samples and cause overfitting. The other method uses transfer learning-based multi-modal representation learning, which transfers the representation of one modality to another (e.g., [122]). The second is the approach for unaligned instances. In this case, the training instances of different modalities are not necessary from the same entities. The first method uses transfer learning to achieve coordinated multimodal representations, e.g., using text messages to improve visual representations for image classification [146]. The second is conceptual grounding. It learns semantic meanings or concepts based on data of multiple modalities, including language, vision, and sound, by finding a common latent space of all multi-modal representations (e.g., [147]). The last one is zero-shot learning, which recognizes a concept without explicitly seeing any examples of it [32]. For example, it can use a neural network to map the text-based message with the extracted visual features. The third is for the hybrid case. In this case, approaches use a pivot modality to learn coordinated multi-modal representations using e.g., bridge correlational neural network, or rely on the existence of large datasets from a similar or related task [32].